# Reality’s Canvas, Language’s Brush: Crafting 3D Avatars from Monocular Video

### 08 Dec 2023 v1

@@ 4.1 Implementation Details
We adopt Adam optimizer [24]. We set the learning rate to $5\times 10^{-4}$ in the initial stage with 200K iterations to warm up the training, where we apply the regularization loss with pseudo labels extracted by pre-trained DensePose [11] network to constrain the learning of UV map and semantic information. After the warm-up stage, we will remove the regularization loss and jointly train all modules with additional mask loss for the rest 200K iterations with $1\times 10^{-4}$ as the learning rate. The whole training stage takes about 60 hours on a single V100 GPU. During inference, our pipeline renders 3.6s per frame with 512 $\times$ 512 image resolution. Hyper-parameters setting and other additional implementation details can be found in the supplementary.
Figure 5: Qualitative results for text instructed human rendering on ZJU-MoCap dataset [45] for novel views. Our approach is capable of generating a consistent human body based on different text instructions. More specifically, our approach enables fine-grained manipulation with the specific module, i.e., shape editing with UVS module, re-lighting with our shading branch, $H_{c}^{N}$, and in the end, texture and appearance editing with our neural texture module.
@@ 4.2 Dataset and Metrics
We train and evaluate our approach on the ZJU-MoCap dataset [45] and People Snapshot dataset [2]. The ZJU-Mocap dataset is captured with 23 hardware-synchronized cameras. For an equal comparison with HumanNeRF [60], we use the same experimental setting by splitting video frames captured by the first camera into seen and unseen pose sets in a 4:1 ratio and use other camera views with the same seen pose as the novel view evaluation set. We repeat this procedure on 6 different subjects and then average the quantitative results. For the People Snapshot dataset, we run EasyMocap [1] to extract the approximate camera and body pose, and then follow the same training and testing split.
@@ 4.3 Comparison with State-of-the-art Methods
Table 1: Quantitative comparision on ZJU-Mocap dataset [45].
@@
Comparison on Novel View Synthesis
Comparison on Novel Pose Rendering
We compare our method against the baselines on novel pose rendering. As listed in Tab. 1, our method shows the best performance for the novel pose scenario with a large margin over UV Volumes [8]. We further present a qualitative comparison between our method and HumanNeRF [60] in Fig. 3, where the neural avatar is driven by novel poses generated by text instruction with the help of MDM [57]. We observe that our method can render the human avatar with better quality and fewer artifacts. Similarly, qualitative results on the People Snapshot dataset [2] in Fig. 4 illustrate that ReCaLaB avoids floating artifacts and animates the neural human avatar in high quality.
Qualitative Results for Texture Editing
ReCaLaB enables editing neural humans with text instructions. For example, with text instructions like "Turn him into a very fat man", "Make the illumination very dim", "Turn his T-shirt red", and "Turn him into a clown", our approach connects the text prompt with the corresponding module and updates the editable neural texture to generate the manipulated neural avatar, as shown in Fig. 5. We also form a human avatar manipulation baseline by combining HumanNeRF [60] and InstructNeRF2NeRF [12]. It is important to note that a HumanNeRF-based manipulation baseline suffers both from a considerably longer 16x training time and lower performance, due to a less-efficient patch-based training scheme. We present the visualization comparison in the supplementary due to short of space.
Figure 6: Qualitative results for human rendering on ZJU-MoCap dataset [45]: The images are in the order of learning neural texture with a) albedo branch $H_{c}^{A}$ only, b) without disentanglement of albedo and shading, c) our full pipeline $H_{c}^{A}\cdot H_{c}^{N}$, and d) Ground Truth. Our approach has achieved the best recognition of the light situation at the calf.
4.4 Ablation Studies
Ablation on Neural Texture Field
We first give an ablation study on the neural texture field, where we investigate the effect of our module on modeling illumination for more photo-realistic rendering. We explore several alternative ways to our approach where we disentangle albedo and shading, including learning neural texture with 1) albedo branch only, and 2) without disentangling the albedo and shading branch. Upon comparison with the alternative methods, our approach stands out in effectively representing the lighting conditions, as highlighted within the red circle in Fig. 6. This improvement can be attributed to our ability to proficiently learn the canonical texture while simultaneously modeling the albedo and shading information to simulate different lighting effects based on varied view directions and human poses. Quantitatively, our full model improves by 8.14% over the ablated version on LPIPS, and more details can be found in the supplementary.
Ablating Current Multi-View Baseline
To showcase the generalization ability of ReCaLaB for novel human pose rendering, we additionally include a comparison with the UVS-map driven multi-view model UV Volumes [8] where all views are used for training. UVV-full denotes a training of [8] with all 20 views, while the same model is trained with a single view in UVV-mono. Tab. 2 provides the quantitative evaluation between our method ReCaLaB and these two baselines. The results demonstrate that our method is on par with UVV-full for novel view synthesis performance. For novel pose rendering, our approach even surpasses the multi-view approach and improves PSNR by approx. 6dB, even though being trained with 19x less data and no synchronized videos. These results show the effectiveness of our method by leveraging human pose priors to map geometry and texture information into a canonical neural texture field.
@@ 4.4 Ablation Studies
For the avatar manipulation task, we study how the neural texture field can influence the editing performance for a given text instruction, e.g., "Make the illumination very dim". We free the parameters of different parts of our neural texture field to make it editable while keeping the other parts frozen. The text-prompted manipulations are shown in Fig. 7. It becomes evident that the shading branch has the most significant influence on the requested illumination change by the text prompt. As we can see, making the shading branch editable satisfies most with regard to the illumination editing requested in the language description, explaining that the shading branch encodes the most illumination-related information and are capable of editing on illumination for neural human.
Figure 7: Qualitative results for illumination change text prompt. The effect of the text instruction "Make the illumination very dim" is shown by updating only selected parts of ReCaLaB, i.e., albedo branch only, shading branch only, and full neural texture module.
@@ 5 Conclusion
In this paper, we propose ReCaLaB, a novel approach designed to acquire an editable neural texture field for human body rendering. Our method offers an efficient methodology of disentangled neural texture learning which can be further edited through textual instructions, making it particularly suitable for monocular video applications. ReCaLaB excels in learning a canonical disentangled neural texture field, yielding exceptional performance in human rendering for both novel view and novel pose scenarios overcoming even recent multi-view setups with just a monocular video input that can be easily captured by everyone. Our differential avatar manipulation pipeline further democratizes human avatar manipulation through a simple language-based interface. We firmly believe that ReCaLaB represents a substantial step forward in the domain of 3D avatar creation and manipulation.

### 24 Mar 2024 v2

@@ 4.1 Implementation Details
We adopt Adam optimizer [28]. We set the learning rate to $5\times 10^{-4}$ in the initial stage with 200K iterations to warm up the training, where we apply the regularization loss with pseudo labels extracted by the pre-trained DensePose [12] network to constrain the learning of the UV map and semantic information. After the warm-up stage, we will remove the regularization loss and jointly train all modules with additional mask loss for the remaining 200K iterations with $1\times 10^{-4}$ as the learning rate. The whole training stage takes about 60 hours on a single V100 GPU. During inference, our pipeline renders 3.6s per frame with 512 $\times$ 512 image resolution. Hyper-parameters setting and other additional implementation details can be found in the supplementary.
@@ 4.2 Dataset and Metrics
We train and evaluate our approach on the ZJU-MoCap dataset [53] and the People Snapshot dataset [2]. The ZJU-Mocap dataset is captured with 23 hardware-synchronized cameras. For an equal comparison with HumanNeRF [70], we use the same experimental setting by splitting video frames captured by the first camera into seen and unseen pose sets in a 4:1 ratio and use other camera views with the same seen pose as the novel view evaluation set. We repeat this procedure on 6 different subjects and then average the quantitative results. For the People Snapshot dataset, we run EasyMocap [1] to extract the approximate camera and body pose, and then follow the same training and testing split.
Figure 4: Qualitative results for human rendering on ZJU-MoCap dataset [53] for novel pose. Our method yields the most superior texture outcomes.
Figure 4: Qualitative results for human rendering on ZJU-MoCap dataset [53] for novel pose. Our method yields the most superior texture outcomes.
Figure 5: Qualitative results on novel poses guided by text instruction. ReCaLaB provides sharper texture and fewer geometric artifacts compared to HumanNeRF [70].
@@ 4.3 Comparison with State-of-the-art Methods
Comparison on Novel View Synthesis
Table 1: Quantitative results for comparing to baselines on ZJU-Mocap dataset [53].
@@
Comparison on Novel Pose Rendering
We compare our method against the baselines on novel pose rendering. As listed in Tab. 2, our method shows the best performance for the novel pose scenario with a large margin over UV Volumes [9], and Fig. 5 indicates that our method provides better texture and fewer geometric artifacts compared to other baselines.
We further present a qualitative comparison between our method and HumanNeRF [70] in Fig. 5, where the neural avatar is driven by novel poses generated by text instruction with the help of MDM [66], a) and b) are frames from the instructions “A person turns to his right and paces back and forth”, and “A person is running” respectively. We observe that our method can render the human avatar with better quality and fewer artifacts. Similarly, qualitative results on the People Snapshot dataset [2] in Fig. 6 illustrate that ReCaLaB avoids floating artifacts and animates the neural human avatar in high quality. Multi-view methods such as UV Volumes [9] heavily depend on input from multiple synchronized cameras which hinders their ability to achieve compelling rendering results in single-view scenarios for novel poses beyond the learned pose domain. We observed that UV Volumes [9] fails to produce adequate results in this challenging setup.
Qualitative Results for Texture Editing
ReCaLaB enables editing neural humans with text instructions. For example, with text instructions like “Turn him into a very fat man”, “Make the illumination very dim”, “Turn his T-shirt red”, and “Turn him into a clown”, our approach connects the text prompt with the corresponding module and updates the editable neural texture to generate the manipulated neural avatar, as shown in Fig. 7.
Figure 7: Qualitative results for text instructed human rendering on ZJU-MoCap dataset [53] for novel views. Our approach is capable of generating a consistent human body based on different text instructions. More specifically, our approach enables fine-grained manipulation with the specific module, i.e., shape editing with UVS module, re-lighting with lighting correction branch, $H_{c}^{N}$, and in the end, texture and appearance editing with our neural texture module.
Comparisons for Texture Editing
We form human avatar manipulation baselines by combining HumanNeRF [70] and UV Volumes [9] with InstructNeRF2NeRF [13], and then update the canonical neural rendering module according to the text instruction. Fig. 8 shows the comparison of the editing performance under the novel pose scenario. Specifically, the reconstructed human avatars from all the methods are edited by the text “turn him into a clown”, and rendered under a novel pose which is triggered by the reference frame. As we can see from Fig. 8, our method provides visually higher quality results with fewer artifacts compared to other baselines, as highlighted with the red circles. It is also important to note that a HumanNeRF-based manipulation baseline suffers both from a considerably longer 16x training time and lower performance, due to a less-efficient patch-based training scheme.
Figure 8: Qualitative results on texture editing. Our method excels in producing the best edited human avatars compared to alternative approaches.
4.4 Ablation Studies
Ablation on Neural Texture Field
We first give an ablation study on the neural texture field, where we investigate the effect of our module on modeling illumination for more photo-realistic rendering. We explore several alternative ways to our approach where we disentangle diffused color and lighting correction branches, including learning neural texture with 1) diffused color branch only (Without-correction), and 2) without disentangling the diffused color and lighting correction branches (Mixture). Upon comparison with the alternative methods, our approach stands out in effectively representing the lighting conditions, as highlighted within the red circle in Fig. 9. This improvement can be attributed to our ability to proficiently learn the canonical texture while simultaneously modeling the diffused color and the lighting correction information to simulate different lighting effects based on varied view directions and human poses.
We further include the detailed quantitative comparison in Tab. 3 under novel pose synthesis for three random subjects, where our full model can achieve the best PSNR over all the alternatives, demonstrating that our approach can model the texture as well as the illumination effectively. Specifically, both our full model and the ablated version with the diffused color branch can generate sharper texture over the baseline model without disentanglement, reflected by the LPIPS metric.
Figure 9: Qualitative results for ablation on Neural Texture Field. Our approach has achieved the best recognition of the light situation at the calf.
Ablating Current Multi-View Baseline
To showcase the generalization ability of ReCaLaB for novel human pose rendering, we additionally include a comparison with the UVS-map driven multi-view model UV Volumes [9] where all views are used for training. UVV-full denotes a training of [9] with all 20 views, while the same model is trained with a single view in UVV-mono. Tab. 2 provides the quantitative evaluation between our method ReCaLaB and these two baselines by evaluating on the same held-out view (View 5). The results demonstrate that our method is on par with UVV-full for novel view synthesis performance. For novel pose rendering, our approach even surpasses the multi-view approach and improves PSNR by approximately 6dB, even though being trained with 19x less data and no synchronized videos. These results show the effectiveness of our method by leveraging human pose priors to map geometry and texture information into a canonical neural texture field.
@@ 4.4 Ablation Studies
For the avatar manipulation task, we study how the neural texture field can influence the editing performance for a given text instruction, e.g., “Make the illumination very dim”. We free the parameters of different parts of our neural texture field to make it editable while keeping the other parts frozen. The text-prompted manipulations are shown in Fig. 3. It becomes evident that the lighting correction branch has the most significant influence on the requested illumination change by the text prompt. As we can see, making the lighting correction branch editable satisfies most with regard to the illumination editing requested in the language description, explaining that the lighting correction branch encodes the most illumination-related information and is capable of editing on illumination for neural humans.
Table 3: Quantitative results for comparing to the ablation on neural texture field under novel pose scenario.
Novel Pose
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Without-correction 29.77 0.972 0.024
Mixture 30.03 0.972 0.029
Ours 30.05 0.972 0.024
Figure 10: Qualitative results for illumination change text prompt. effect of the text instruction “Make the illumination very dim” is shown by updating only selected parts of ReCaLaB.
@@ 5 Conclusion
In this paper, we propose ReCaLaB, a novel approach designed to acquire an editable neural texture field for human body rendering. Our method offers an efficient methodology of disentangled neural texture learning which can be further edited through textual instructions, making it particularly suitable for monocular video applications. ReCaLaB excels in learning a disentangled neural texture field, yielding exceptional performance in human rendering for both novel view and novel pose scenarios overcoming even recent multi-view setups with just a monocular video input that can be easily captured by everyone. Our differential avatar manipulation pipeline further democratizes human avatar manipulation through a simple language-based interface. We firmly believe that ReCaLaB represents a substantial step forward in the domain of 3D avatar creation and manipulation.
Appendix S1 Additional Experimental Results
Quantitative results for each subject
As aforementioned in the main paper, we compared our approach and state-of-the-art methods for neural human rendering on 6 subjects in the ZJU-MoCap dataset [53], including 313, 377, 387, 392, 393, and 394. Tab. S1 and Tab. S2 show the detailed quantitative results of the novel view and novel pose scenarios for all subjects. As we can see from the quantitative results, our approach largely outperforms the UV Volumes [9] on all the evaluation metrics. Meanwhile, compared to HumanNeRF [70], our approach has shown competitive performance for novel view scenarios, while performing better on novel pose synthesis for all subjects. The quantitative comparisons demonstrate the effectiveness of our approach for neural human reconstruction by learning the canonical neural texture and disentangling the shape and appearance information.
Qualitative results for novel view
In Fig. 3 of the main paper, we have shown the qualitative results for three subjects. We further show the qualitative results of all the rest subjects in Fig. S1. HumanNeRF [70] generates disconnected body texture with dark lines (see marked regions), and the UV Volumes [9] method suffers from low rendering quality. In contrast, our approach successfully addresses both issues and provides results visually closer to the ground truth.
Subject 313 Subject 377 Subject 387
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
UV Volumes [9] 24.59 0.965 0.043 23.53 0.969 0.041 21.75 0.957 0.049
HumanNeRF [70] 29.57 0.967 0.030 30.35 0.975 0.024 27.94 0.962 0.037
Ours 29.40 0.968 0.029 30.40 0.977 0.021 27.73 0.962 0.037
Subject 392 Subject 393 Subject 394
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
UV Volumes [9] 24.22 0.965 0.047 22.54 0.956 0.053 23.24 0.958 0.047
HumanNeRF [70] 30.51 0.969 0.033 28.53 0.961 0.037 29.85 0.961 0.036
Ours 30.54 0.970 0.033 28.34 0.961 0.038 29.69 0.962 0.036
Table S1: Novel view synthesis quantitative comparison on ZJU-MoCap dataset. We show the results of each subject.
Subject 313 Subject 377 Subject 387
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
UV Volumes [9] 23.36 0.968 0.043 23.56 0.979 0.035 21.97 0.962 0.045
HumanNeRF [70] 29.85 0.970 0.027 31.02 0.980 0.019 27.87 0.963 0.035
Ours 30.36 0.972 0.026 30.98 0.981 0.016 27.49 0.963 0.035
Subject 392 Subject 393 Subject 394
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
UV Volumes [9] 25.27 0.974 0.035 22.20 0.961 0.046 22.02 0.961 0.043
HumanNeRF [70] 31.77 0.975 0.026 28.84 0.965 0.033 28.37 0.962 0.032
Ours 31.55 0.975 0.026 28.75 0.965 0.032 28.82 0.963 0.031
Table S2: Novel pose synthesis quantitative comparison on ZJU-MoCap dataset. We show the results of each subject.
Figure S1: Qualitative results for human rendering on ZJU-MoCap dataset [53] for novel view synthesis. Our approach gives the best quality for the rendered human body.
Text Editing Qualitative Assessment for Diverse Subjects.
We present our results on additional subjects in Figure S2, which suggests the versatility of our approach, and showcases its applicability across different subjects.
Appendix S2 Additional Implementation Details
As described in the main manuscript, our approach reconstructs the human avatar with the disentangled neural feature and enables different manipulations including novel view rendering, novel pose synthesis, and human avatar editing with language brush. During the training stage, we first employ the photo-metric loss $\mathcal{L}_{\text{rec}}$ to obtain the densities and colors along the bent rays using the deformation module to establish accurate deformation learning in the initial stage, which is similar to HumanNeRF. We then learn the UVS map and neural texture with the alternative training strategy to optimize regularization loss $\mathcal{L}_{\text{reg}}$ from Eq. 10 in the main paper and the MSE loss while freezing the deformation module. Then we use $\mathcal{L}_{\text{rec}}$ with a learning rate $1\times 10^{-4}$ to optimize the rendering details by training UVS map and neural texture jointly. In the end, we further introduce the mask loss and jointly train all the modules for enhancing the neural rendering quality and geometric accuracy.
When editing with a language brush, the text instruction will guide our model to update the corresponding modules to efficiently edit the shape, illumination, and appearance, where we iteratively supervise our model with training objectives described in Eq. 9 in the main paper from edited ground-truth to fulfill the instructed editing while maintaining the multi-view geometrical consistency.
Figure S2: Qualitative results for text Editing for novel view synthesis.